### v4

# UniAP: Unifying Inter- and Intra-Layer Automatic Parallelism by Mixed Integer Quadratic Programming

Abstract
Distributed learning is commonly used for training deep learning models, especially large models. In distributed learning, manual parallelism (MP) methods demand considerable human effort and have limited flexibility. Hence, automatic parallelism (AP) methods have recently been proposed for automating the parallel strategy optimization process. Existing AP methods suffer from sub-optimal solutions because they do not jointly optimize the two categories of parallel strategies (i.e., inter-layer parallelism and intra-layer parallelism). In this paper, we propose a novel AP method called UniAP, which unifies inter- and intra-layer automatic parallelism by mixed integer quadratic programming. To the best of our knowledge, UniAP is the first parallel method that can jointly optimize the two categories of parallel strategies to find an optimal solution. Experimental results show that UniAP outperforms state-of-the-art methods by up to 3.80$\times$ in throughput and reduces strategy optimization time by up to 107$\times$ across five Transformer-based models.
1 Introduction
Deep learning models have demonstrated promising performance across many domains. For example, deep learning models such as BERT [1], GPT-3 [2] and T5 [3] achieve state-of-the-art (SOTA) performance on many natural language processing (NLP) tasks. For computer vision (CV), deep learning models such as ViT [4] and Swin Transformer [5] achieve good accuracy on multiple tasks.
Distributed learning (also called parallel learning) on clusters with several machines or GPUs is commonly used for training deep learning models, especially for some large models with billions of parameters [2; 6; 7]. Several parallel strategies, including pipeline parallelism (PP), data parallelism (DP), tensor parallelism (TP), and fully sharded data parallelism (FSDP), have been proposed for distributed learning. These parallel strategies can be divided into two main categories: inter-layer parallelism and intra-layer parallelism. Inter-layer parallelism [8; 9; 10; 11; 12; 13; 14; 15], which includes PP, partitions the model into disjoint sets without partitioning tensors in each layer. Intra-layer parallelism [16; 17; 18; 19], which includes DP, TP, and FSDP, partitions tensors in a layer along one or more axes.
The parallel method [^1] To avoid confusion, we treat ‘parallel method’ and ‘parallel strategy’ as two different terminologies in this paper. in one specific distributed learning method or system typically adopts one parallel strategy or a combination of several parallel strategies. Existing parallel methods can be divided into two categories: manual parallelism (MP) methods and automatic parallelism (AP) methods. In MP methods  [20; 18; 21], one or several parallel strategies are manually optimized by researchers or developers. MP methods require extensive domain knowledge in deep learning models and hardware architectures. With the rapid development of deep learning models and the increasing diversity of modern hardware architectures [22; 23], MP methods demand considerable human effort and have limited flexibility.
To address the two limitations of MP methods, AP methods [9; 24; 25] have recently been proposed for automating the parallel strategy optimization process. Although existing AP methods have achieved promising progress, they optimize the two categories of parallel strategies separately rather than jointly. More specifically, some methods optimize only one category of parallel strategies [26; 27; 28; 29; 30; 31; 32], and the others optimize inter- and intra-layer parallelism hierarchically [9; 33; 34; 11; 24; 25]. Hence, existing AP methods suffer from sub-optimal solutions.
In this paper, we propose a novel AP method called UniAP for distributed learning. The contributions of UniAP are outlined as follows:
  • UniAP unifies inter- and intra-layer automatic parallelism by mixed integer quadratic programming (MIQP) [35].
  • To the best of our knowledge, UniAP is the first parallel method that can optimize the two categories of parallel strategies jointly rather than separately to find an optimal solution.
  • Experimental results show that UniAP outperforms state-of-the-art methods by up to 3.80$\times$ in throughput and reduces strategy optimization time by up to 107$\times$ across five Transformer-based models.
2 Background
2.1 Parallel strategy
Pipeline parallelism (PP) In PP, each worker (machine or GPU) holds a subset of model layers. Adjacent layers on different workers need to transfer activations in the forward propagation (FP) step and gradients in the backward propagation (BP) step.
Data parallelism (DP) In DP, each worker holds a replica of the whole model and partitions training samples. In each iteration, each worker computes gradients and synchronizes them with the other workers using all-reduce collective communication (CC). All workers will have the same model parameters after the synchronization step.
Tensor parallelism (TP) In TP, each worker holds a replica of training samples and partitions within model layers. In each iteration, each worker computes its local outputs in FP and its local gradients in BP. To synchronize outputs and gradients, all workers will perform all-reduce CC in FP and BP steps according to the partition scheme.
Fully sharded data parallelism (FSDP) FSDP partitions optimizer states, parameters and gradients of the model into separate workers. During the FP and BP step of each iteration, FSDP performs an all-gather CC to obtain the complete parameters for the relevant layer, respectively. After computing gradients, FSDP conducts a reduce-scatter CC to distribute the global gradients among the workers.
2.2 Parallel method
Manual parallelism (MP) MP refers to the parallel methods in which human experts design and optimize the parallel strategies. Representative MP methods include Megatron-LM [18], Mesh-TensorFlow [20], and GSPMD [21]. Megatron-LM manually designs TP and PP strategies for training Transformer-based models and exhibits superior efficiency. Mesh-TensorFlow and GSPMD require human effort to designate and tune the intra-layer parallel strategy. These methods rely on expert design and have little flexibility, challenging their automatic application to other models.
Inter-layer-only AP or intra-layer-only AP For inter-layer-only AP, GPipe [8] and vPipe [30] employ a balanced partition algorithm and a dynamic layer partitioning middleware to partition pipelines, respectively. For intra-layer-only AP, OptCNN [26], TensorOpt [31], and Tofu [27] employ dynamic programming methods to optimize DP and TP strategies together. FlexFlow [28] and Automap [29] use the Monte Carlo method to find the optimal DP and TP strategy. Colossal-Auto [32] utilizes integer programming (IP) techniques to generate intra-layer parallelism and activation checkpointing strategies. All these methods optimize only one category of parallel strategies.
Inter- and intra-layer AP PipeDream [9], DAPPLE [11], and PipeTransformer [24] use dynamic programming to determine optimal strategies for both DP and PP. DNN-partitioning [33] adopts IP and dynamic programming to explore DP and PP strategies. DT-FM [36] combines genetic algorithm and the local search strategy to explore DP and PP strategies in a layer-wise manner. Piper [34] and Alpa [25] adopt a parallel method considering DP, TP, and PP. Galvatron [37] uses dynamic programming to determine DP, TP, and FSDP strategies in a single pipeline stage. As for PP, it partitions stages and determines micro-batch size using naive greedy algorithms. All these methods are hierarchical, which will result in sub-optimal solutions.
3 Method
In this section, we introduce our proposed method called UniAP. Apart from current AP methods, UniAP jointly optimizes the two categories of parallel strategies, including PP, DP, TP, and FSDP, to find an optimal solution. Visualization on the differences between existing parallel methods and UniAP is available in Appendix A.
Figure 1: Flowchart of UniAP.
Figure 1 illustrates the flowchart of UniAP. UniAP first profiles the runtime information for the user’s hardware environment and the deep learning model. After that, UniAP estimates inter- and intra-layer costs given the computation graph and profiling results with its cost models. The estimated costs and the computation graph are then transformed into an MIQP problem. The objective function of the MIQP is to maximize the training throughput, or in other words, to minimize the training time per iteration (TPI). By iteratively applying the cost model and MIQP with different parameters, UniAP determines the minimal TPI and its corresponding parallel strategies. We name this process the Unified Optimization Process (UOP). Finally, UniAP interprets the parallel strategies into the execution plan for the designated model.
3.1 Profiling
UniAP collects runtime information about the hardware environment and deep learning model during profiling. For the hardware environment, UniAP evaluates the efficiency of all-reduce and point-to-point (P2P) communication for different device subsets. For example, when profiling a node with 4 GPUs, UniAP measures the all-reduce efficiency for various DP, TP, and FSDP combinations across these GPUs. Additionally, UniAP ranks these GPUs from 0 to 3 and evaluates the speed of P2P for two pipeline options: ($0\rightarrow 2$ and $1\rightarrow 3$) and ($0\rightarrow 1$, $1\rightarrow 2$ and $2\rightarrow 3$). Furthermore, UniAP estimates the computation-communication overlap coefficient (CCOC) [37; 38].
UniAP acquires two types of information for the deep learning model: computation time and memory usage. On one hand, UniAP distinguishes the forward computation time per sample for different types of hidden layers. On the other hand, UniAP collects memory usage information for each layer, including the memory occupied by parameters and the memory usage of activation per sample in different TP sizes.
3.2 Cost model
UniAP employs two primary cost models, namely the time cost model and the memory cost model.
Time cost model To estimate computation time, UniAP first calculates the forward computation time by multiplying the batch size with the forward computation time per sample obtained from profiling. Users can obtain a more precise result by profiling the forward time on the specified batch size. For Transformer-based models that mainly consist of the MatMul operator, the computation time in the BP stages is roughly twice that of the FP stages [18; 12; 37]. Additionally, UniAP estimates the communication time by dividing the size of transmitting tensors by the profiled communication efficiency for different communication primitives. To accommodate overlapping, UniAP multiplies the profiled CCOC by the overlapping interval of computation and communication. To model the communication time between pipeline stages, UniAP calculates the cross-stage cost between consecutive stages by the summation of P2P costs.
Memory cost model UniAP estimates memory consumption for each layer with its memory cost model. This estimation consists of three steps for a given layer. First, it computes the activation memory cost $m_{a}$ by multiplying the batch size and the profiled activation memory cost per sample of the TP size used by the strategy. Next, UniAP calculates the memory cost of model states $m_{s}$ for each layer based on their parameter size $ps$, TP size $ts$, FSDP size $fs$, and a constant $c_{dtype}$ dependent on the data type. Formally, we have
$m_{s}=\frac{c_{dtype}\times ps}{ts\times fs}.$ (1)
For example, if we choose precision as FP32, then $c_{dtype}=(4+4+4+4)/4=4$ since the learnable parameters, gradients, momentum and variance consume equal size of memory. If we opt for mixed precision with FP16 activated, then $c_{dtype}=(4+4+4+2+2)/2=8$. Finally, UniAP aggregates the activation memory cost $m_{a}$, memory cost of model states $m_{s}$, and context memory cost $m_{c}$ to a constant matrix ${\boldsymbol{\mathit{M}}}$, where ${\boldsymbol{\mathit{M}}}_{uk}$ denotes the memory cost for the $k$-th intra-layer strategy of layer $u$ on a single device.
3.3 Mixed integer quadratic programming
The estimated costs and the computation graph are then transformed into an MIQP problem. Its formulation includes an objective function and several constraints.
3.3.1 Objective function
Figure 2: Time cost decomposition of a GPipe-style PP.
Figure 3: A contiguous set.
The objective function tries to minimize TPI. In this paper, we have chosen GPipe as our PP strategy for illustration. [^2] UniAP is also compatible with other PP strategies. For example, users need to modify only the memory constraint in Section 3.3.2 to adapt to synchronous 1F1B pipeline [10; 11]. Figure 3 depicts the time cost decomposition of a GPipe-style PP with non-negligible communication costs. The time needed to apply gradients at the end of each iteration is not included, as it depends on the optimizer and is insignificant compared to the total time spent on FP and BP.
We denote the cost for computation stages as =$\mathbb{P}=\{p_{1},p_{2},\dots,p_{pp\_size}\}$ and the cost for communication stages as $\mathbb{O}=\{o_{1},o_{2},\dots,o_{pp\_size-1}\}$. Here, $pp\_size$ represents the number of computation stages, which corresponds to the pipeline parallel size. In Figure 3, $fp_{i}$ and $bp_{i}$ denote forward and backward computation time for computation stage $i$, respectively. $fo_{j}$ and $bo_{j}$ denote forward and backward communication time for communication stage $j$, respectively. Hence, we have $p_{i}=fp_{i}+bp_{i}$ and $o_{j}=fo_{j}+bo_{j}$.
In a GPipe-style pipeline, we use $c$ to denote the number of micro-batches. As illustrated in Figure 3, a mini-batch is uniformly split into four micro-batches, and the total TPI is determined by the latency of all computation and communication stages and the latency of the slowest stage. We further denote TPI in GPipe as $tpi_{gpipe}$. Given that a stage with a higher FP computation cost leads to a higher BP computation cost with high probability, we can write the objective function of GPipe-style pipeline as follows:
$\min tpi_{gpipe}=\sum_{i=1}^{pp\_size}p_{i}+\sum_{j=1}^{pp\_size-1}o_{j}+(c-1)\max\left(\mathbb{P}\cup\mathbb{O}\right).$ (2)
3.3.2 Constraint
We first introduce additional notations before presenting the constraints. For a given layer $u\in{\mathbb{V}}$, ${\mathbb{S}}_{u}$ represents its set of intra-layer parallel strategies, ${\boldsymbol{\mathit{A}}}_{uk}$ denotes the $k$-th intra-layer execution cost obtained from our time cost model. Additionally, we use ${\boldsymbol{\mathit{S}}}_{uk}\in\{0,1\}$ to indicate whether the $k$-th parallel strategy is selected for the layer $u$, and use ${\boldsymbol{\mathit{P}}}_{ui}\in\{0,1\}$ to indicate whether layer $u$ is placed on the $i$-th computation stage. Each edge $\langle u,v\rangle\in{\mathbb{E}}$ is assigned a resharding cost denoted by ${\boldsymbol{\mathit{R}}}_{uv}$ if the vertices are located within the same pipeline stage. Alternatively, if the vertices are located across consecutive stages, the resharding cost between them is denoted by ${\boldsymbol{\mathit{R}}}^{\prime}_{uv}$. These two resharding costs are constant matrices derived from our time cost model.
Computation-stage constraint To compute the total cost for a single computation stage $i$, all computation and communication costs associated with that stage must be aggregated and assigned to $p_{i}$. This constraint can be formulated as follows:
$\sum_{u\in{\mathbb{V}}}{\boldsymbol{\mathit{P}}}_{ui}{\boldsymbol{\mathit{S}}}_{u}^{\mathsf{T}}{\boldsymbol{\mathit{A}}}_{u}+\sum_{\langle u,v\rangle\in{\mathbb{E}}}{\boldsymbol{\mathit{P}}}_{ui}{\boldsymbol{\mathit{P}}}_{vi}({\boldsymbol{\mathit{S}}}_{u}^{\mathsf{T}}{\boldsymbol{\mathit{R}}}_{uv}{\boldsymbol{\mathit{S}}}_{v})=p_{i},~{}\forall i\in\{1,\dots,pp\_size\}.$ (3)
On the left side of Equation (3), the first polynomial term represents the cost of choosing specific intra-layer strategies for layers placed in stage $i$. The second term represents total resharding costs within stage $i$.
Communication-stage constraint To calculate the total cost for a single communication stage $j$, we should aggregate the P2P costs incurred between consecutive stages and assign them to $o_{j}$. This constraint can be formulated as follows:
$\sum_{\langle u,v\rangle\in{\mathbb{E}}}{\boldsymbol{\mathit{P}}}_{uj}{\boldsymbol{\mathit{P}}}_{v(j+1)}({\boldsymbol{\mathit{S}}}_{u}^{\mathsf{T}}{\boldsymbol{\mathit{R}}}^{\prime}_{uv}{\boldsymbol{\mathit{S}}}_{v})$ $=o_{j},~{}\forall j\in\{1,\dots,pp\_size-1\}.$ (4)
Memory constraint We need to guarantee that no devices (GPUs) will encounter out-of-memory (OOM) exceptions during training process. This constraint can be formulated as follows:
$\sum_{u\in{\mathbb{V}}}{\boldsymbol{\mathit{P}}}_{ui}{\boldsymbol{\mathit{S}}}_{u}^{\mathsf{T}}{\boldsymbol{\mathit{M}}}_{u}\leqslant m,~{}\forall i\in\{1,\dots,pp\_size\}.$ (5)
Here, $m$ denotes the memory limit for each device. In the case of homogeneous computing devices, the value of $m$ remains constant throughout all stages. But the value of $m$ varies in the case of heterogeneous computing devices.
Order-preserving constraint PP is not a single-program multiple-data (SPMD) parallel strategy [8]. Hence, we need an order-preserving constraint to ensure that the subgraphs of $\mathcal{G}$ are contiguous. We adopt the definition of contiguous from previous work [33; 34].
Definition 3.1.
A set ${\mathbb{W}}\subseteq{\mathbb{V}}$ is contiguous if there do not exist nodes $u\in{\mathbb{W}}$, $v\in{\mathbb{V}}\setminus{\mathbb{W}}$, and $w\in{\mathbb{W}}$ such that $v$ is reachable from $u$ and $w$ is reachable from $v$.
Figure 3 illustrates an example of a contiguous set ${\mathbb{W}}$, in which we cannot find any reachable node pairs $\langle u,v\rangle$ and $\langle v,w\rangle$ where $u,w\in{\mathbb{W}}$ and $v\in{\mathbb{V}}\setminus{\mathbb{W}}$.
In our case, our model will not be assigned to different pipeline stages in a disordered manner if we ensure that all subgraphs on each computation stage are contiguous. After reformulating this constraint in linear form, we have
${\boldsymbol{\mathit{Z}}}_{vi}$ $\geqslant{\boldsymbol{\mathit{P}}}_{vi},~{}$ $\forall v\in{\mathbb{V}},~{}\forall i\in\{1,2,\dots,pp\_size\},$ (6a)
${\boldsymbol{\mathit{Z}}}_{vi}$ $\leqslant{\boldsymbol{\mathit{Z}}}_{ui},~{}$ $\forall u,v\in{\mathbb{V}},~{}\forall\langle u,v\rangle\in{\mathbb{E}},~{}\forall i\in\{1,2,\dots,pp\_size\},$ (6b)
${\boldsymbol{\mathit{Z}}}_{vi}$ $\leqslant{\boldsymbol{\mathit{P}}}_{vi}-{\boldsymbol{\mathit{P}}}_{ui}+1,~{}$ $\forall u,v\in{\mathbb{V}},~{}\forall\langle u,v\rangle\in{\mathbb{E}},~{}\forall i\in\{1,2,\dots,pp\_size\}.$ (6c)
Detailed proof can be found in Appendix B.
Layer-placement constraint All layers should be placed on exactly one pipeline stage and at least one layer should be placed on each pipeline stage. This constraint can be formulated as follows:
$\sum_{i=1}^{pp\_size}{\boldsymbol{\mathit{P}}}_{ui}=1,$ $\forall u\in{\mathbb{V}},$ (7a)
$\sum_{u\in{\mathbb{V}}}{\boldsymbol{\mathit{P}}}_{ui}\geqslant 1,$ $\forall i\in\{1,\dots,pp\_size\},$ (7b)
${\boldsymbol{\mathit{P}}}_{ui}\in\{0,1\},$ $\forall u\in{\mathbb{V}},~{}i\in\{1,\dots,pp\_size\}.$ (7c)
Strategy-selection constraint Each layer must and can choose only one strategy. This constraint can be formulated as follows:
$\sum_{k=1}^{\lvert{\mathbb{S}}_{u}\rvert}{\boldsymbol{\mathit{S}}}_{uk}=1,$ $\forall u\in{\mathbb{V}},$ (8a)
${\boldsymbol{\mathit{S}}}_{uk}\in\{0,1\},$ $\forall u\in{\mathbb{V}},~{}k\in\{1,\dots,|{\mathbb{S}}_{u}|\}.$ (8b)
The MIQP formulation for UniAP includes the objective function in Equation (2) and all the constraints from Euqation (3) - (8b).
3.4 Unified optimization process
Algorithm 1 Unified Optimization Process
  Input: Profiling results $PR$, strategy dictionary $SD$, mini-batch size $B$, computation graph $\mathcal{G}$, and the number of GPUs $n$.
  Output: Optimal cost $cost^{*}$, pipeline parallel size $pp\_size^{*}$, the number of micro-batches $c^{*}$, layer placement ${\boldsymbol{\mathit{P}}}^{*}$, and intra-layer strategy ${\boldsymbol{\mathit{S}}}^{*}$
  $pp\_size^{*}=1$;
  $c^{*}=B$;
  $A$, $R$, _, $M$ = CostModeling($PR$, $SD[1]$, $\mathcal{G}$, $B$);
  $cost^{*}$, $P^{*}$, $S^{*}$ = QIP(${\boldsymbol{\mathit{A}}}$, ${\boldsymbol{\mathit{R}}}$, ${\boldsymbol{\mathit{M}}}$);
  Get all factors for $n$ except 1 and insert them to set ${\mathbb{F}}$;
  Get all factors for $B$ except 1 and insert them to set ${\mathbb{B}}$;
  for $pp\_size$ in ${\mathbb{F}}$ do
     for $c$ in ${\mathbb{B}}$ do
        Micro-batch size $b=B/c$;
        ${\boldsymbol{\mathit{A}}}$, ${\boldsymbol{\mathit{R}}}$, ${\boldsymbol{\mathit{R}}}^{\prime}$, ${\boldsymbol{\mathit{M}}}$ = CostModeling($PR$, $SD[pp\_size]$, $\mathcal{G}$, $b$);
        $cost$, ${\boldsymbol{\mathit{P}}}$, ${\boldsymbol{\mathit{S}}}$ = MIQP(${\boldsymbol{\mathit{A}}}$, ${\boldsymbol{\mathit{R}}}$, ${\boldsymbol{\mathit{R}}}^{\prime}$, ${\boldsymbol{\mathit{M}}}$, $pp\_size$, $c$);
        if $cost<cost^{*}$ then
           $cost^{*}$, $pp\_size^{*}$, $c^{*}$, ${\boldsymbol{\mathit{P}}}^{*}$, ${\boldsymbol{\mathit{S}}}^{*}$ = $cost$, $pp\_size$, $c$, ${\boldsymbol{\mathit{P}}}$, ${\boldsymbol{\mathit{S}}}$;
        end if
     end for
  end for
UOP integrates the cost model and MIQP based on the profiling results and the computation graph to return the optimal parallel strategy and the corresponding TPI. Algorithm 1 summarizes the whole process. In Algorithm 1, we denote intra-layer cost as ${\boldsymbol{\mathit{A}}}$, inter-layer cost as ${\boldsymbol{\mathit{R}}}$, cross-stage cost as ${\boldsymbol{\mathit{R}}}^{\prime}$, and memory cost as ${\boldsymbol{\mathit{M}}}$. The CostModeling process calculates these four costs based on the cost model described in Section 3.2.
First, UOP optimizes intra-layer-only parallelism for cases in which pipeline parallelism is not adopted. Several works [25; 32] have used quadratic integer programming (QIP) to optimize intra-layer-only parallel strategy and achieved promising results. UniAP provides a QIP formulation for intra-layer-only parallelism in Appendix C.
Then, UOP enumerates all factors of $n$ except 1 as the pipeline parallel size $pp\_size$. For each $pp\_size$, UOP enumerates all factors of $B$ except 1 as the number of micro-batches $c$. These enumerations aim to achieve load balance on a homogeneous cluster.
For each candidate $pp\_size$ and $c$, UOP formulates the cost for a training iteration to an MIQP expression. It then waits for the MIQP solver to return the optimal cost and parallel strategy under the current configuration. Specifically, our implementation adopts the Gurobi Optimizer [39] as our QIP and MIQP solver.
Finally, UOP returns the minimum cost $cost^{*}$ and its corresponding pipeline parallel size $pp\_size^{*}$, number of micro-batches $c^{*}$, layer placement ${\boldsymbol{\mathit{P}}}^{*}$, and intra-layer strategies ${\boldsymbol{\mathit{S}}}^{*}$. We provide visualization for a candidate solution to UOP in Appendix D.
3.5 Complexity analysis
Let $\lvert{\mathbb{V}}\rvert$, $\lvert{\mathbb{S}}\rvert$, and $n$ denote the number of layers, parallel strategies, and GPUs, respectively. As illustrated in Algorithm 1, UniAP enumerates all factors of $n$ except 1 as the $pp\_size$ in the outer loop and all factors of $B$ except 1 in the inner loop. Since the time complexity of enumerating factors of any positive integer $a~{}(a\in{\mathbb{N}}_{+})$ is $\mathcal{O}(\sqrt{a})$, the time complexity for the two loops in the UniAP algorithm is $\mathcal{O}(\sqrt{Bd})$. Within the inner loop body, UniAP calls CostModeling to model the cost of each stage for each parallel strategy. Furthermore, the optimization time limit of the MIQP solver can be set as a constant hyperparameter when UniAP calls it. Therefore, the overall computational complexity of UniAP is $\mathcal{O}(\lvert{\mathbb{V}}\rvert\lvert{\mathbb{S}}\rvert\sqrt{Bd})$.
4 Experiment
We conduct experiments on five different kinds of environments. EnvA refers to a node (machine) with 1 Xeon 6248 CPU, 8 V100-SXM2 32GB GPUs, and 472GB memory. EnvB refers to 2 nodes interconnected with 10Gbps networks, where each node has 2 Xeon E5-2620 v4 CPUs, 4 TITAN Xp 12GB GPUs, and 125GB memory. EnvC refers to a node with 8 A100 40GB PCIe GPUs. EnvD has 4 nodes, each with the same configuration as that in EnvB. EnvE is a cloud cluster with 8 nodes and each node has 4 DCUs (a type of non-NVIDIA GPU).
We evaluate UniAP with five Transformer-based models, BERT-Huge [1], T5-Large [3], ViT-Huge [4], Swin-Huge [5], and Llama [6; 7]. We follow the common practice of training these transformer-based models. To eliminate factors that affect training throughput, we turn off techniques orthogonal to parallel strategies, such as activation checkpointing [40]. However, we integrate FP16 mixed precision training [41] for the largest model, Llama, to successfully orchestrate the process. More details on these models are provided in Appendix E.
The experimental evaluation concentrates on two primary metrics: training throughput and strategy optimization time. The former is calculated by averaging throughput from the 10th to the 60th iteration of training, while the latter is determined by measuring the time of the UOP. More details are provided in Appendix E.
Table 1: Training throughput and strategy optimization time on five Transformer-based models.
Env. Model Training throughput (samples/s) Minimum speedup Maximum speedup
Galvatron Alpa UniAP
EnvA BERT-Huge 33.46 $\pm$ 0.28 31.56 $\pm$ 0.04 33.46 $\pm$ 0.28 1.00 1.06
T5-Large 23.29 $\pm$ 0.04 MEM×$\times$2) 23.29 $\pm$ 0.04 1.00 1.00
ViT-Huge 109.51 $\pm$ 0.07 97.66 $\pm$ 1.42 109.51 $\pm$ 0.07 1.00 1.12
Swin-Huge CUDA×$\times$3) N/A4) 67.96 $\pm$ 0.12 N/A4) N/A4)
EnvB BERT-Huge 6.27 $\pm$ 0.17 8.95 $\pm$ 0.06 10.77 $\pm$ 0.13 1.20 1.71
T5-Large1) 8.06 $\pm$ 0.06 MEM×$\times$2) 7.98 $\pm$ 0.05 0.99 0.99
ViT-Huge 32.20 $\pm$ 0.17 38.74 $\pm$ 0.20 45.58 $\pm$ 0.54 1.18 1.41
Swin-Huge 13.90 $\pm$ 0.17 N/A4) 19.08 $\pm$ 0.10 1.37 1.37
EnvC Llama-7B 1.22 $\pm$ 0.01 N/A4) 4.63 $\pm$ 0.007 3.80 3.80
Env. Model Strategy optimization time (min.) Minimum speedup Maximum speedup
Galvatron Alpa UniAP
EnvA BERT-Huge 6.44 $\pm$ 0.588 $>$ 40 0.37 $\pm$ 0.002 17.29 $>$ 107.41
T5-Large 12.41 $\pm$ 0.122 MEM×$\times$2) 0.89 $\pm$ 0.007 13.98 13.98
ViT-Huge 6.29 $\pm$ 0.464 $>$ 40 0.57 $\pm$ 0.009 10.95 $>$ 69.60
Swin-Huge 11.88 $\pm$ 0.666 N/A4) 2.16 $\pm$ 0.004 5.49 5.49
EnvB BERT-Huge 2.04 $\pm$ 0.010 $>$ 40 1.51 $\pm$ 0.005 1.34 $>$ 26.32
T5-Large1) 2.64 $\pm$ 0.110 MEM×$\times$2) 0.91 $\pm$ 0.005 2.90 2.90
ViT-Huge 2.37 $\pm$ 0.180 $>$ 40 1.11 $\pm$ 0.011 2.14 $>$ 36.01
Swin-Huge 4.29 $\pm$ 0.320 N/A4) 2.29 $\pm$ 0.010 1.87 1.87
EnvC Llama-7B 6.84 $\pm$ 0.055 N/A4) 0.58 $\pm$ 0.006 11.83 11.83
  • 1)
    T5-Large tested on EnvB is restricted to 16/16 layers to avoid out-of-memory (OOM) exceptions.
  • 2)
    MEM×$\times$: OOM exceptions during strategy optimization.
  • 3)
    CUDA×$\times$: CUDA OOM exceptions during model training.
  • 4)
    The official implementation for Swin-Huge and Llama in Alpa is absent. We have endeavored to adopt the code, yet encountered compilation errors. Consequently, experiments related to these models on Alpa are marked as N/A.
4.1 Training throughput and strategy optimization time
We compare the training throughput and strategy optimization time of UniAP with those of the baselines on EnvA, EnvB, and EnvC. For experiments conducted on EnvA, we set the mini-batch size to be 32, 16, 128, and 128 for BERT, T5, ViT, and Swin, respectively. For experiments conducted on EnvB, we set the mini-batch size to be 16, 8, 64, and 32 for these four models, respectively. For the Llama model [6; 7] run on EnvC, we set the mini-batch size to be 8.
We choose Galvatron [37] and Alpa [25] as baselines because they have achieved SOTA performance. Specifically, Galvatron has surpassed other methods, including PyTorch DDP [16], Megatron-LM [18], FSDP [42; 19], GPipe [8], and DeepSpeed 3D [43] in terms of training throughput, as reported in the original paper [37]. Furthermore, Alpa utilizes the Just-In-Time (JIT) compilation feature in JAX and outperforms Megatron-LM and DeepSpeed.
Table 1 shows the training throughput and strategy optimization time on EnvA, EnvB, and EnvC. On EnvA, UniAP and Galvatron get the same optimal strategy for BERT-Huge, T5-Large, and ViT-Huge, outperforming Alpa in terms of training throughput and strategy optimization time. In addition, UniAP finds a solution for Swin-Huge, while Galvatron encounters CUDA OOM problems. In particular, UniAP achieves a maximum optimization speedup that is 17$\times$ faster than Galvatron and hundreds of times faster than Alpa on BERT-Huge. This is mainly due to the ability of the MIQP solver to search for an optimal strategy on multiple threads, while the dynamic programming based methods like Galvatron and Alpa run on a single thread due to their strong data dependency.
On EnvB, UniAP consistently demonstrates competitive or larger training throughput compared to Galvatron and Alpa. We attribute the performance improvement to the larger strategy space in UniAP. A detailed study of the importance of strategy space is provided in Section 4.4. Furthermore, the strategy optimization time of UniAP is also significantly shorter than the two baseline methods.
On EnvC, UniAP shows an optimization speedup of 11.83$\times$ and a training speedup of 3.80$\times$ compared to Galvatron on Llama. We examine the parallel strategy they adopted. The results show that UniAP employs an 8-stage PP with a micro-batch size of 1, whereas Galvatron employs a 4-stage PP with a micro-batch size of 3 (with the final micro-batch containing 2 samples). Within each PP stage, Galvatron utilizes a 2-way TP. For EnvC, which is equipped with 8 A100 40GB PCIe GPUs, minimizing communication volume is critical. Since PP has significantly less inter-device communication volume than TP, the parallel strategy discovered by UniAP is more reasonable than Galvatron in this case.
To facilitate further discussions, we provide another case study of the optimal parallel strategy for BERT-Huge in Appendix F.
4.2 Estimation accuracy
Some variables in UniAP and other AP methods are estimated values rather than actual running values. The TPI (inverse of training throughput) is the most important one. Accurate estimation for TPI or training throughput is crucial for evaluating candidate parallel strategies and ensuring the optimality of the solution. We hereby introduce a metric called relative estimation error (REE) $e$ to quantify the accuracy of the estimated training throughput $\hat{T}$ to the actual training throughput $T$:
$e(T,\hat{T})=\frac{|T-\hat{T}|}{T}\times 100\%.$ (9)
We evaluate the optimal parallel strategies obtained from EnvA and EnvB. Results show that UniAP achieves an average REE of 3.59%, which is lower than 11.17% for Galvatron.
4.3 Scalability
(a) Training throughput.
(b) Strategy optimization time.
Figure 4: Scalability on training throughput and strategy optimization time.
We study the scalability of UniAP using BERT, T5, ViT, and Swin on EnvD. We set the mini-batch sizes for each data point as 8, 4, 32, and 16 times the number of nodes (denoted as ‘#nodes’). The experimental results are shown in Figure 4. In Figure 4(a), the training throughput of the optimal strategy demonstrate near-linearity as the number of nodes and mini-batch size increase. In Figure 4(b), the strategy optimization time matches the complexity analysis in Section 3.5.
Additionally, we use EnvE to scale UniAP to larger clusters and larger models including Llama-13B. UniAP also outperforms other baselines in this experiment. Details are available in Appendix G.
4.4 Ablation study
Table 2: Ablation study on the importance of unifying strategy space.
Model Training throughput (samples/s)
UniAP (Inter-only) UniAP (Intra-only) UniAP
BERT-Huge SOL×$\times$1) 2.48 $\pm$ 0.02 10.77 $\pm$ 0.13
T5-Large SOL×$\times$1 2.92 $\pm$ 0.01 9.01 $\pm$ 0.06
ViT-Huge 45.58 $\pm$ 0.54 CUDA×$\times$2 45.58 $\pm$ 0.54
Swin-Huge 19.08 $\pm$ 0.10 4.66 $\pm$ 0.02 19.08 $\pm$ 0.10
  • 1)
    SOL×$\times$: No solution after strategy optimization.
  • 2)
    CUDA×$\times$: CUDA OOM exceptions during model training.
We investigate the importance of the strategy space for the optimality of parallel strategies with an ablation study. Specifically, we constrain the strategy space to inter-layer-only and intra-layer-only strategies and evaluate the training throughput of the resulting optimal strategy on EnvB. We set the mini-batch sizes to be 16, 12, 64, and 32, respectively.
Results are shown in Table 2. We can find that constraining the strategy space compromises the optimality of parallel strategies or gets strategies that encounter OOM across different models. Hence, unifying inter- and intra-layer AP for joint optimization is essential and necessary.
5 Conclusion
In this paper, we propose a novel AP method called UniAP to unify inter- and intra-layer AP by MIQP. To the best of our knowledge, UniAP is the first parallel method that can jointly optimize the two categories of parallel strategies to find an optimal solution. Experimental results show that UniAP can outperform other state-of-the-art baselines to achieve the best performance.
Appendix
Appendix A Visualization for parallel methods
In order to provide an intuitive analysis, we present visualizations in Figure 5 on different parallel methods for optimizing parallel strategies for a three-layer model. The numbered boxes represent the corresponding layers of the model. The layers with gray backgrounds represent the final parallel strategy for each method. The layers without dotted lines indicate that DP was applied to them, while TP was applied to the others. The direction of the dotted lines indicates the dimension of the tensor that TP is applied to. The solid arrows between layers represent connections within a single pipeline stage, while the dashed arrows represent connections between consecutive pipeline stages. For example, the parallel strategy optimized by MP (Figure 5(b)) represents that PP and TP are utilized to train the model. Specifically, layer 0 of the model is assigned to the first pipeline stage, while layer 1 and 2 are assigned to the second pipeline stage. Additionally, TP is activated across all model layers.
As shown in Figure 5(b), MP methods are manually designed and rely on expert knowledge, resulting in limited flexibility. Inter-layer-only and intra-layer-only AP methods (Figure 5(c) and 5(d)) optimize (search) from a set of candidate inter-layer-only and intra-layer-only parallel strategies, respectively. To the best of our knowledge, existing inter- and intra-layer parallelism AP methods optimize parallel strategies hierarchically. Thus, we refer to them as Hierarchical AP in Figure 5(e). They initially adopt algorithms, such as greedy or dynamic programming, to propose candidate inter-layer parallel strategies. Subsequently, they optimize the intra-layer parallel strategy for every fixed inter-layer parallel strategy. Compared to them, UniAP jointly optimize all possible parallel strategies from intra-layer-only parallelism to inter- and intra-layer parallelism. Consequently, UniAP has the largest strategy space for exploration.
Appendix B Proof of the linear form for the contiguous set
To facilitate our discussion, we adopt the linear form of the order-preserving constraint as presented in the main paper. We denote ${\boldsymbol{\mathit{P}}}_{ui}$ as a 0-1 variable indicating whether layer $u$ is to be placed on the $i$-th computation stage, $pp\_size$ as the number of computation stages in the pipeline. Besides, $\mathcal{G}({\mathbb{V}},{\mathbb{E}})$ represents the computation graph for the model. Then, we formalize the theorem as follows:
Theorem B.1.
A subgraph ${\mathbb{V}}_{i}=\{\forall u\in{\mathbb{V}}:{\boldsymbol{\mathit{P}}}_{ui}=1\}$ is contiguous if and only if there exists ${\boldsymbol{\mathit{Z}}}_{vi}$ such that Equation (6a), (6b), and (6c) are satisfied.
Previous work [33] has proven this theorem. Our proof draws on the process of this work. The details of the proof are as follows:
Proof.
"If": Assume that there exists nodes $u,w\in{\mathbb{V}}_{i}$ and $v\notin{\mathbb{V}}_{i}$ such that $v$ and $w$ are reachable from $u$ and $v$, respectively. Hence, ${\boldsymbol{\mathit{P}}}_{ui}=1$, ${\boldsymbol{\mathit{P}}}_{wi}=1$, and ${\boldsymbol{\mathit{P}}}_{vi}=0$. Without losing generality, we assume $\langle u,v\rangle\in{\mathbb{E}}$. Thus, according to Equation (6c), we have ${\boldsymbol{\mathit{Z}}}_{vi}\leqslant{\boldsymbol{\mathit{P}}}_{vi}-{\boldsymbol{\mathit{P}}}_{ui}+1=0$. By applying Equation (6b) repeatedly following the path from $v$ to $w$, we have ${\boldsymbol{\mathit{Z}}}_{wi}\leqslant{\boldsymbol{\mathit{Z}}}_{vi}$. Thus, ${\boldsymbol{\mathit{Z}}}_{wi}\leqslant 0$. However, we also have ${\boldsymbol{\mathit{Z}}}_{wi}\geqslant{\boldsymbol{\mathit{P}}}_{wi}=1$ according to Equation (6a). A contradiction.
"Only if": First, we define ${\boldsymbol{\mathit{Z}}}_{vi}=1$ if a node $w\in{\mathbb{V}}_{i}$ is reachable from $v~{}(v\in{\mathbb{V}})$. Otherwise, ${\boldsymbol{\mathit{Z}}}_{vi}=0$. Thus, Equation (6a) and (6b) are satisfied according to this kind of definition. For Equation (6c), if ${\boldsymbol{\mathit{P}}}_{vi}=1$, the constraint will hold true regardless of whether ${\boldsymbol{\mathit{P}}}_{ui}$ is $1$ or 00. If ${\boldsymbol{\mathit{P}}}_{vi}=0$ and ${\boldsymbol{\mathit{P}}}_{ui}=0$, ${\boldsymbol{\mathit{Z}}}_{vi}\leqslant{\boldsymbol{\mathit{P}}}_{vi}-{\boldsymbol{\mathit{P}}}_{ui}+1=1$ will also hold true because ${\boldsymbol{\mathit{Z}}}_{vi}$ could be either 00 or $1$. Finally, if ${\boldsymbol{\mathit{P}}}_{vi}=0$ and ${\boldsymbol{\mathit{P}}}_{ui}=1$, ${\boldsymbol{\mathit{Z}}}_{vi}=0$ will hold true because ${\mathbb{V}}_{i}$ is a contiguous set and we cannot find any $w\in{\mathbb{V}}_{i}$, such that $w$ is reachable from $v$. ∎
(a) Model
(b) MP
(c) Inter-layer-only AP
(d) Intra-layer-only AP
(e) Hierarchical AP
(f) UniAP
Figure 5: Parallel methods for optimizing parallel strategies for a three-layer model in (a).
Appendix C QIP formulation for intra-layer-only parallelism
Here we present the QIP formulation for intra-layer-only parallelism with explanations.
Objective function In terms of intra-layer-only parallelism, there is only one computation stage involved. As a result, the objective function takes into account only the value of $p_{1}$. We hereby formalize the equation as
$\min\quad tpi_{gpipe}=p_{1}.$ (10)
Computation-stage constraint With only one computation stage in intra-layer-only parallelism, the communication-stage constraint can be omitted, and the computation and communication cost can be modeled for $p_{1}$. Thus, we could formalize the constraint as
$\sum_{u\in{\mathbb{V}}}{\boldsymbol{\mathit{S}}}_{u}^{\mathsf{T}}{\boldsymbol{\mathit{A}}}_{u}+\sum_{\langle u,v\rangle\in{\mathbb{E}}}{\boldsymbol{\mathit{S}}}_{u}^{\mathsf{T}}{\boldsymbol{\mathit{R}}}_{uv}{\boldsymbol{\mathit{S}}}_{v}=p_{1}.~{}$ (11)
In the equation, the first summation term for any $u\in{\mathbb{V}}$ represents the cost of choosing intra-layer strategies for all layers, while the second term represents the summation of resharding costs on all edges.
Memory constraint Similar to the memory constraint in MIQP, it is necessary to ensure that the memory usage on a single device does not exceed its device memory bound $m$ in QIP. This restriction gives
$\sum_{u\in V}{\boldsymbol{\mathit{S}}}_{u}^{\mathsf{T}}{\boldsymbol{\mathit{M}}}_{u}\leqslant m.~{}$ (12)
It is worth noting that $m$ should be an identical constant across multiple devices if these devices are homogeneous. Otherwise, the value of $m$ varies.
Strategy-selection constraint For intra-layer-only parallelism, the layer-placement constraint can be safely omitted because it is designed for PP. However, the strategy-selection constraint is necessary because each layer can only select one intra-layer strategy. Therefore, the strategy-selection constraint for QIP is identical to Equation (8a) and (8b) for MIQP.
By combining objective function (10) and constraints (8a), (8b), (11), and (12), we have the QIP expression for optimizing the intra-layer-only AP. Like MIQP expression for optimizing the inter- and intra-layer AP, UniAP will eventually get the minimum TPI and corresponding parallel strategies by invoking the off-the-shelf solver.
Appendix D Visualization for the candidate solution
Figure 6: A candidate solution for UOP.
In this section, we proceed to visually represent a potential solution for UOP. Given a deep learning model $\mathcal{G}$, pipeline parallel size $pp\_size$, and number of micro-batches $c$, UniAP will determine layer placement ${\boldsymbol{\mathit{P}}}$ for inter-layer parallelism and parallel strategy ${\boldsymbol{\mathit{S}}}$ for intra-layer parallelism using an off-the-shelf solver. As Figure 6 shows, the solver is optimizing a three-layer model with two pipeline stages, each assigned four GPUs. At this time, a candidate solution could be
${\boldsymbol{\mathit{P}}}=\begin{bmatrix}1&0\\ 1&0\\ 0&1\end{bmatrix},~{}{\boldsymbol{\mathit{S}}}=\begin{bmatrix}1&0&0\\ 0&1&1\\ 0&0&0\\ \vdots&\vdots&\vdots\\ 0&0&0\end{bmatrix}.$ (13)
Here, the $u$-th row of matrix ${\boldsymbol{\mathit{P}}}$ denotes the placement strategy for layer $u$, where ${\boldsymbol{\mathit{P}}}_{ui}=1$ signifies the placement of layer $u$ on stage $i$, while 00 indicates otherwise. For example, ${\boldsymbol{\mathit{P}}}_{l_{0}}=\left[1,~{}0\right]$ denotes the placement of layer $l_{0}$ on pipeline stage 1. Additionally, the $u$-th column of matrix ${\boldsymbol{\mathit{S}}}$ denotes the selected intra-layer parallel strategy for layer $u$, where ${\boldsymbol{\mathit{S}}}_{uj}=1$ denotes the selection of the $j$-th strategy from the intra-layer parallel strategy set. For example, ${\boldsymbol{\mathit{S}}}_{l_{0}}=\left[1,~{}0,~{}0,~{}\cdots,~{}0\right]^{\mathsf{T}}$ indicates that layer $l_{0}$ will adopt only the DP strategy, while ${\boldsymbol{\mathit{S}}}_{l_{1}}=\left[0,~{}1,~{}0,~{}\cdots,~{}0\right]^{\mathsf{T}}$ indicates that layer $l_{1}$ will employ a strategy where DP is performed on GPU 0, 1 and GPU 2, 3, and TP is performed across these two GPU groups.
There exist numerous combinations of ${\boldsymbol{\mathit{P}}}$ and ${\boldsymbol{\mathit{S}}}$. The off-the-shelf solver will automatically search for the optimal solution given pipeline parallel size $pp\_size$ and the number of micro-batches $c$. By solving the MIQP expression and enumerating every possible $pp\_size$ and $c$ in the UOP process, UniAP will ultimately derive an optimal parallel strategy for the deep learning model within the current hardware environment.
Appendix E Experiment detail
Gurobi configuration When tackling the MIQP problem, UniAP employs several configurations for the Gurobi Optimizer 10.1 [39]. In particular, we set TimeLimit to 60 seconds, MIPFocus to 1, NumericFocus to 1, and remain other configurations to default. For instance, we establish the MIPGap parameter as the default value of 1e-4 to serve as a strict termination criterion. Furthermore, we have implemented an early stopping mechanism to terminate the optimization process as early as possible. There are two conditions that can activate the mechanism. Firstly, if the current runtime exceeds 15 seconds and the relative MIP optimality gap is less than 4%, we will terminate the optimization. Secondly, if the current runtime exceeds 5 seconds and the best objective bound is worse than the optimal solution obtained in the previous optimization process, we will terminate the optimization.
Table 3: Summary of the evaluated models.
Model Task1) Statistics Precision
#hidden layers Hidden size Sequence length #params
BERT-Huge PT 32 1280 512 672M FP32
T5-Large CG 24/24 1024 512 737M FP32
ViT-Huge IC 32 1280 196 632M FP32
Swin-Huge IC 2/2/42/2 320 49 $\times$ 64 1.02B FP32
Llama-7B CLM 32 4096 2048 7B FP16
Llama-13B CLM 40 5120 2048 13B FP16
  • 1)
    PT: Pretraining; CG: Conditional Generation; IC: Image Classification; CLM: Causal Language Modeling.
Model detail Table 3 summarizes six Transformer-based models selected for our evaluations. Four of these models, namely BERT-Huge [1], T5-Large [3], Llama-7B, and Llama-13B [6, 7], belong to the domain of natural language processing (NLP). At the same time, the remaining two, ViT-Huge [4] and Swin-Huge [5], are associated with computer vision (CV). It is noteworthy that BERT, ViT, and Llama maintain consistent types of hidden layers respectively, whereas T5 and Swin have different types of hidden layers. Numbers separated by slashes represent the statistical information for different layer types. For instance, Swin-Huge comprises four types of layers, each with 2, 2, 42, and 2 layers, respectively.
Training detail UniAP is based on the PyTorch framework and integrates models from HuggingFace Transformers. It employs various types of parallelism, including Pipeline Parallelism (PP), Data Parallelism (DP), Tensor Parallelism (TP), and Fully Sharded Data Parallelism (FSDP), utilizing GPipe [8], PyTorch DDP [16], Megatron-LM [18], and FairScale [19], respectively. For NLP models, we use the English Wikipedia dataset [44], while the ImageNet-1K dataset [45] is used for CV models. We train these models using the Adam optimizer [46]. We omit hyperparameters here such as learning rate and weight decay as these have minimal impact on training throughput. The model parameters in the HuggingFace Transformers are configured to align with the specifications of each individual model. For instance, we set hidden_size to 1280, num_hidden_layers to 32, num_attention_heads to 16, and seq_length to 512 for BERT-Huge. Regarding other hyperparameters in the HuggingFace configurations, we set hidden_dropout_prob and attention_probs_dropout_prob to 0.0 for ViT-Huge. For Swin-Huge, we set drop_path_rate to 0.2. We remain other configurations to default. It should be noted that the training batch sizes for each experiment are outlined in the main paper.
Appendix F Case study: BERT-Huge
Figure 7: The optimal parallel strategy for all hidden layers of BERT-Huge on EnvB. Different colors represent different input samples in a micro-batch.
In this section, we present a visualization of the optimal parallel strategy discovered by UniAP. As represented in Figure 7, the strategy pertains to training BERT-Huge with 32 hidden layers in a 2-node environment EnvB with a mini-batch size of 16. Each node was equipped with 2 Xeon E5-2620 v4 CPUs, 4 TITAN Xp 12GB GPUs, and 125GB memory. These nodes are interconnected via a 10Gbps network. It should be noted that we only showcase the parallel strategy for the hidden layers here for simplicity but without losing generality.
Here, we provide further topological information for a node within EnvB. As illustrated in Figure 8, we categorize the GPUs numbered 0 and 1 in each node and refer to them collectively as GPUGroup0. Similarly, we label the GPUs numbered 2 and 3 as GPUGroup1. In EnvB, the interconnects within each GPU group (i.e., PCIe) have superior bandwidth than that between different groups (i.e., QPI). We collectively designate these two connection bandwidths as intra-node bandwidth, which is higher than inter-node bandwidth.
Figure 8: Topology of a node in EnvB.
In this example, UniAP has identified a parallel strategy for inter-layer parallelism that involves a two-stage pipeline. This strategy utilizes parallelism in a manner that is both efficient and effective. Specifically, the communication cost of point-to-point (P2P) between two nodes is less than that of all-reduce. Given that the inter-node bandwidth is lower than the intra-node bandwidth, the two-stage PP becomes a reasonable choice. Moreover, the pipeline has been designed such that each stage comprises an equal number of layers. This design leverages the homogeneity of the nodes and ensures load balancing across the cluster.
Within each PP stage, UniAP employs an intra-layer parallel strategy. It utilizes a 2-way DP for the initial 12 hidden layers in each stage between GPUGroup0 and GPUGroup1. For the remaining four hidden layers, a 2-way FSDP is utilized between GPUGroup0 and GPUGroup1 to reduce memory footprint and meet memory constraints. Within each GPU group, UniAP employs a 2-way TP for each layer. In general, TP incurs more significant communication volumes than DP and FSDP. In order to achieve maximum training throughput on EnvB, it is necessary to implement parallel strategies that prioritize higher communication volumes within each group and lower volumes between groups. Therefore, the strategy for BERT-Huge with 32 hidden layers combines the best elements of PP, DP, TP, and FSDP to maximize training throughput.
In addition, we have conducted calculations for the model FLOPs utilizatio (MFU)[47] for Galvatron, Alpa, and UniAP in this scenario to validate our analysis. MFU is independent of hardware, frameworks, or implementations. Therefore, it allows us to examine the performance of different parallel strategies solely from a strategic perspective. For BERT-Huge, the resulting MFUs for UniAP, Galvatron, and Alpa are 58.44%, 58.44%, and 55.10% on EnvA, while 23.6%, 13.7%, and 19.6% on EnvB, respectively. These results validate that UniAP’s optimization of inter- and intra-layer AP results in superior performance compared to Galvatron and Alpa.
Appendix G Scaling to larger clusters and larger models
We adapt UniAP to a cloud cluster to further scale UniAP to larger clusters and larger models. The experiments are conducted on 8 nodes. Each of them has 1 Intel Xeon Phi Processor 7285, 4 16GB PCIe DCU, 128GB DDR4 DRAM, and 200Gb Infiniband interconnection network. We name this environment as EnvE. It is worth noting that DCU is a kind of GPU, which is different from NVIDIA GPUs. We select MP methods including Megatron [18] and DeepSpeed (ZeRO-3) [17] as our baseline methods.
In the experiment, we evaluate these baselines and UniAP with Llama-7B and Llama-13B [6]. We set the mini-batch size as 8 for Llama-7B and 4 for Llama-13B. The remaining configurations align with those for Llama-7B in Section 4, such as activating FP16 mixed precision training.
We focus on two metrics: training throughput and strategy optimization time. For UniAP, the former metric is calculated by averaging throughput from the 10th to the 60th iteration of training, while the later metric is determined by measuring the time of the UOP (Section 3.4). Please note that Megatron [18] and DeepSpeed [17] do not automatically optimize parallel strategies. Consequently, we undertake an exhaustive approach, examining every combination of parallel strategies and the corresponding parallel configurations they accommodate. For each combination, we assess the average throughput over the 10th to 60th training iteration, ultimately adopting the fastest strategy as their optimization result. The total time spent in the process is recorded as their strategy optimization time.
Table 4: Training throughput and strategy optimization time on EnvE.
Model Training throughput (samples/s) Strategy optimization time (min.)
Megatron DeepSpeed UniAP Megatron DeepSpeed UniAP
Llama-7B 2.01 $\pm$ 0.005 SOL×$\times$1) 2.01 $\pm$ 0.005 > 8.0 hours SOL×$\times$1) 3.07 $\pm$ 0.121
Llama-13B 0.82 $\pm$ 0.001 SOL×$\times$1) 0.82 $\pm$ 0.001 > 2.5 hours SOL×$\times$1) 1.95 $\pm$ 0.076
  • 1)
    SOL×$\times$: No solution after strategy optimization.
As shown in Table 4, UniAP consistently identifies the fastest parallel strategies while expending considerably less time on strategy optimization compared to Megatron (about 157$\times$ faster for Llama-7B and 77$\times$ faster for Llama-13B). Please note that the strategy optimization time Megatron needs for Llama-7B surpasses that for Llama-13B. We attribute this phenomenon to the variations in the mini-batch size. Specifically, the escalation from a mini-batch size of 4 (employed in Llama-13B) to 8 (employed in Llama-7B) leads to a rise in the number of candidate parallel strategies, as well as the parallel strategies that will successfully train the model. As a result, the strategy optimization time will become increasingly long when the mini-batch size exceeds 8 in the pretraining scenario [2, 6, 7]. In this situation, UniAP will identify the optimal parallel strategy much faster.
Furthermore, our experiments highlight a limitation encountered with DeepSpeed (ZeRO-3). It requires the mini-batch size to be divisible evenly by the total number of computation devices. This specific prerequisite prevents DeepSpeed from successfully launching the training process with 32 DCUs.
Table 5: Statistics on the candidate parallel strategies for Megatron.
Model Batch size Training throughput (samples/s) #infeasible5) #candidate6)
Top-11) Top-22) Slowest3) Median4)
Llama-7B 8 2.01 1.92 0.22 0.82 41 64
Llama-13B 4 0.82 0.58 0.27 0.42 42 48
  • 1)
    Top-1: The training throughput achieved by the fastest parallel strategy.
  • 2)
    Top-2: The training throughput achieved by the second fastest parallel strategy.
  • 3)
    Slowest: The training throughput achieved by the slowest parallel strategy.
  • 4)
    Median: The median value of training throughputs across parallel strategies that will successfully train the model.
  • 5)
    #infeasible: The number of parallel strategies that will encounter exceptions such as CUDA OOM during model training.
  • 6)
    #candidate: The total number of candidate parallel strategies.
We further examine the statistics on the candidate parallel strategies for Megatron, as shown in Table 5. When the users are unable to identify the fastest parallel strategy out of hundreds of candidates, they may be required to sacrifice at least 4.4% of the training throughput for Llama-7B and 29.2% for Llama-13B. Even if they successfully eliminate all infeasible parallel strategies, which constitute the majority of candidates (Table 5), they still only have a 50% chance of identifying a parallel strategy with training throughput better than 0.82 samples/s for Llama-7B and 0.42 samples/s for Llama-13B. Consequently, it is challenging for users to identify the fastest parallel strategy for larger clusters and larger models. In such circumstances, UniAP is able to demonstrate its capacity to scale and accommodate both larger clusters and larger models.